# Information Theory Strikes Back: New Development in the Theory of Cardinality Estimation

### v1

@@ 1 Introduction
In this paper, we discuss upper bounds on the query output size that use more refined data statistics: the $\ell_{p}$-norms of degree sequences of the join attributes in the relations. They were originally introduced in our prior work [15].
@@
For more background on cardinality estimation, we point the reader to a recent overview on traditional and pessimistic estimators [14], a recent survey on learned cardinality estimators [9], and a recent monograph on query optimizers, with an discussion on the cardinality estimation in Microsoft SQL Server [8].
@@ 2 Preliminaries
We first introduce the class of queries and data statistics under consideration and then give a brief background in information theory.

### v2

@@ 1 Introduction
In this paper, we overview the upper bounds on the query output size that were previously introduced in our prior work [18]. They use more refined data statistics: the $\ell_{p}$-norms of degree sequences of the join attributes in the relations.
@@
Our upper bounds are not the first to use degree sequences and $\ell_{p}$-norms. Jayaraman et al. [16] introduced an algorithm for evaluating a full conjunctive query and proved a runtime (and implicitly the query output size) in terms of $\ell_{p}$-norms on degree sequences. Their result is limited to binary relations, to a single value $p$ for a given query, and to queries whose length of the minimal cycle is $\geq p+1$. Their upper bound is a special case of our $\ell_{p}$-bounds [1]. The Degree Sequence Bound (DSB) [8] is a tight upper bound of a query $Q$ in terms of the degree sequences of its join attributes. The query $Q$ is restricted to be full Berge-acyclic conjunctive query. DSB and our framework are incomparable [1].
For more background on cardinality estimation, we point the reader to an overview on traditional and pessimistic estimators [17], a survey on learned cardinality estimators [11], and a monograph on query optimizers [10].
@@ 2 Preliminaries
We introduce the class of queries and data statistics under consideration and some background in information theory.